OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -36 and 2304, respectively.

OpenStudy (anonymous):

an = 9 • 4^n an = 9 • 4^(n - ) an = 9 • (-4)^)n + 1) an = 9 • (-4)^(n - 1)

OpenStudy (anonymous):

Do you know the formula for a geometric series?

OpenStudy (anonymous):

sorry sequence?

OpenStudy (anonymous):

I do not

OpenStudy (anonymous):

What is it?@peachpi

OpenStudy (anonymous):

\[a _{n}=a _{1}r ^{n-1}\] a1 = first term and r = common ratio

OpenStudy (anonymous):

since we have the second term we can write \[-36=a _{1}r ^{2-1}=a _{1}r\]

OpenStudy (anonymous):

Can you write the equation for the 5th term?

OpenStudy (anonymous):

2304 = a1r^2 = a1r ?

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

Since it's the 5th term, the exponent is 5-1 = 4

OpenStudy (anonymous):

2304 = a1 * r^4

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So now you have two equations and two unknowns. You can use substitution to solve the system

OpenStudy (anonymous):

-36 = a1*r r = 36/a1 Substitute into the 2304 equation